### v2

# The Schwurbelarchiv: a German Language Telegram dataset for the Study of Conspiracy Theories

Abstract
Sociality borne by language, as is the predominant digital trace on text-based social media platforms, harbours the raw material for exploring multiple social phenomena. Distinctively, the messaging service Telegram provides functionalities that allow for socially interactive as well as one-to-many communication. Our Telegram dataset contains over 6,000 groups and channels, 40 million text messages, and over 3 million transcribed audio files, originating from a data-hoarding initiative named the “Schwurbelarchiv” (from German schwurbeln: speaking nonsense). This dataset publication details the structure, scope, and methodological specifics of the Schwurbelarchiv, emphasising its relevance for further research on the German-language conspiracy theory discourse. We validate its predominantly German origin by linguistic and temporal markers and situate it within the context of similar datasets. We describe process and extent of the transcription of multimedia files. Thanks to this effort the dataset uniquely supports multimodal analysis of online social dynamics and content dissemination. Researchers can employ this resource to explore societal dynamics in misinformation, political extremism, opinion adaptation, and social network structures on Telegram. The Schwurbelarchiv thus offers unprecedented opportunities for investigations into digital communication and its societal implications.
Keywords telegram dataset, conspiracy theories, social media, misinformation
1 Introduction
The proliferation of digital trace data from social media platforms has enriched the study of societal phenomena, enabling researchers to analyze communication patterns and the dissemination of information at large scales Freelon (2014). Platforms like Telegram, with its blend of interactive groups and broadcast channels, serve as fertile grounds for examining issues ranging from political discourse to the spread of conspiracy theories. The insights gained from such analyses reveal their influence on societal narratives.
The Schwurbelarchiv dataset contributes to a growing number of published Telegram datasets for research, e.g. La Morgia et al. (2023); Baumgartner et al. (2020), and – with the transcription of audio files – offers a unique starting point to widen the scope of research into conspiracy theories. This publication aims to provide a resource for researchers to enable large-scale investigations of the societal and linguistic dynamics that are connected with conspiracy theories, misinformation, political dissatisfaction, and extremism online, which has become a growing research interest in recent years Urman and Katz (2022); Curley et al. (2022); Bovet and Grindrod (2022); Hoseini et al. (2023); Willaert et al. (2022); Zehring and Domahidi (2023). The following paper first gives insights into scope and creation process of the dataset with an emphasis on transcribed audio data, and an estimation of how much of the German-language conspiracy-related discourse on Telegram is covered in our dataset.
1.0.1 Telegram.
Telegram is an instant messaging platform where users can exchange messages by creating an account with their phone number. Users can communicate in direct chats (one-to-one communication), in groups of up to 200,000 users (many-to-many), or participate in broadcast channels (one-to-many). These broadcast channels allow an unlimited amount of Telegram users to follow a single Telegram user’s (the administrator’s) content, and therefore broadcast channels are well designed for information dissemination at a large scale. While one-to-one communication is private and generally not accessible to researchers, groups and channels are public and accessible if their URL is known. We refer to groups and channels together as “chats” in the remainder of this work.
2 Related Works
A number of studies on Telegram datasets or research primarily using Telegram data have been conducted in recent years, many of which focus on research on the dissemination of misinformation and conspiracy theories in online environments. Here, we introduce the three Telegram datasets most closely related to our work in terms of temporal and linguistic scope of the data contained. We analyse similarities and differences across datasets, aiming to provide an assessment of the coverage of the Schwurbelarchiv and situating the dataset within the relevant research context.
2.0.1 Public Telegram Datasets.
The study by  Zehring and Domahidi (2023) investigates the connections between the Querdenken movement, a prominent anti-COVID protest network in German-speaking countries often linked to conspiracy theories, to far-right and alternative media, focusing on content-sharing patterns and the influence of QAnon within the network. Using a dataset of 6,294,955 text messages from 578 German public Telegram channels, the research explores the movement’s key topics over time and across its sub-communities. The dataset was collected by first identifying seed channels by searching for actors connected to the Querdenken and far-right movements, to the esoteric/spirituality scene as well as nominees for the Golden Tin Foil Hat prize  [^1] https://www.dergoldenealuhut.de in Germany, Austria, and Switzerland. This approach resulted in 238 initial channels, from which all messages were scraped. Next, frequently forwarded channels (mentioned 50 times or more) were identified and included in the dataset. This procedure – also called “snowball sampling” – was performed twice and led to a total of 578 public Telegram channels from which all messages available for the time between October 28, 2015 and January 3, 2022 were collected. By performing topic modelling on the dataset, Zehring and Domahidi (2023) found four overarching topics present in the Telegram channels: promotion, QAnon, right-wing populism, and COVID-19 conspiracy theories.
A larger (at the time of writing unpublished) dataset on similar topics was assembled by Mohr (2023). Since snowball sampling on its own quickly yields an unmanageable amount of chats, they implemented a guidance mechanism to prioritize chats likely to yield relevant data by predicting their relevance based on a keyword list which was translated into 44 languages. The keywords and therefore selection of chats was targeted toward content containing misinformation related to COVID-19. The dataset which is to our knowledge one of the largest Telegram datasets available, includes 433,000 chats, with full data from 128,000 chats comprising 2.1 billion messages, primarily in Russian, English, and German. Messages, spanning text, metadata, and links (but excluding videos, images, and voice messages), predominantly reflect activity from European time zones and trends linked to events like the European COVID-19 vaccination campaign and the start of the Russian invasion of Ukraine in the time period of January 2020 to January 2023. All predominantly German language chats (23,000) from this dataset were made available to the authors of this study for comparison to the Schwurbelarchiv.
The dataset collected by Baumgartner et al. (2020), the largest publicly available collection of Telegram data to date, includes 27,000 chats and 317 million messages. The dataset is available for the years 2015-2019 and was regularly updated during that time, with source code available to facilitate further data collection, enabling researchers to extract information from specific chats. Starting with a seed list of 250 English-language chats – 124 focused on right-wing extremism and 137 on cryptocurrency – the number of chats included was also expanded using a snowball sampling approach. All of the studies described here used the Telegram API to build customized clients and collect data.
2.0.2 Message Deletion.
The deletion of messages by users on Telegram has emerged as a significant factor influencing the completeness and reliability of Telegram datasets. The study by Buehling (2024) examines the phenomenon of message deletion in Telegram chats, focusing on its implications for computational social science. Using a repeatedly scraped dataset of messages from 25 groups and 25 channels, the study highlights how the nature of digital trace data can lead to systematic biases in analyses such as topic modelling, dictionary-based computational content analysis, and network studies. The study finds that for channels, on average, only 88% of messages remain accessible five days after posting, dropping to 83% after seven months. In groups, the decline is even steeper, with only 64% of messages remaining after five days and 52% after seven months, emphasizing the need for timely data collection. Furthermore, Buehling (2024) finds that the types of messages deleted vary substantially between channels and groups. In groups, text messages are disproportionately deleted, while in channels, forwarded content and videos are more frequently removed. These discrepancies introduce biases not only in computational content analyses but also in the mapping of information flows and actor networks and can skew conclusions drawn about the nature and dynamics of online communication.
3 Methodology
3.0.1 Data Collection.
The entire raw data contained in the Schwurbelarchiv is available on the Internet Archive [^2] https://archive.org/details/schwurbel-archiv [^3] https://web.archive.org/web/20241126233137/https://schwurbelarchiv.wordpress.com/. The person initially responsible for collecting the data remains anonymous, but responds to requests via email and provides information about the data selection process. They describe their motivation for collecting the dataset as the pure joy of data-hoarding, contribution to the Internet Archive, and enablement of research and journalism using the collected data as a source.
Similar to the other Telegram datasets described here, groups and channels in the Schwurbelarchiv were selected via a snowball sampling approach. Starting with a seed set of Telegram chats focused on renowned German speaking conspiracy theorists, the process involved the automated isolation of invitation links to other Telegram chats and joining the chats to collect the messages posted there. Relevance of chats identified this way, e.g. a focus on conspiracy-related content, was determined by a human-in-the-loop to limit the number of included chats. After identification of a relevant chat, its content was initially downloaded in bulk, after which a continual streaming process for content posted in the chat was implemented. In contrast to the data collection approach of the other studies described here, the data was gathered using a Windows Virtual Machine, with four separate Telegram Desktop sessions, each operating under its own account. Chats were then exported using the Remote Desktop Protocol (RDP). A Python script handled the zipping of files to optimize storage, extracted URLs, and uploaded the data. Local files were removed after the upload. RDP enables storing not only the HTML text files, but all video files, voice messages, and images posted in the chats. Video files larger than 8 MB have not been downloaded until mid of 2021, whereafter the limit for downloaded videos was increased to 50 MB. To our knowledge, the Schwurbelarchiv is the only publicly available Telegram dataset that contains multimedia content. This content constitutes a large part of the discourse taking place on social media platforms that is typically ignored due to the difficulty of handling diverse data formats and large data volumes. The fact that our dataset retains this content greatly expands the types of research questions and analysis approaches that it enables.
3.0.2 Data Cleaning and Preprocessing
Since the raw collected data contains multimedia files, it requires a storage space of nearly 24 TB in a compressed state. The sheer size of the data as well as the native Internet Archive file format .warc complicate processing and analysis. Therefore, next to clarifying ethical and licensing questions associated with the use of the Schwurbelarchiv for research (see Section “Ethics and Data Protection” below), we set out to preprocess the dataset to make it more easily accessible for a wider research community. This preprocessing involved data cleaning, transcription of audio content, and the validation of the completeness of the dataset. We describe these steps in more detail in the remainder of this section.
After downloading the raw data from the Internet Archive, we used Python scripts to extract all relevant information, namely author, posting date, and message body. In addition, we distinguish between original messages and forwarded messages – and their respective authors and posting dates – which enables the reconstruction of message forwarding networks within the dataset. Furthermore, we isolate links to websites shared in the message text body. A full listing of all data columns contained in the cleaned dataset is provided in the appendix 5. While the raw dataset as it is published on the Internet Archive contains uncensored usernames, we pseudonymise usernames using a sha-256 hash-function. We extract usernames from the HTML files containing the content of the chats and do not have access to user IDs. As a result, we cannot distinguish between users that have chosen the same username. This is a substantial limitation of our dataset.
During the initial data collection, there was no quality control of the included chats with respect to research interests such as the investigation of conversation patterns. However, the raw dataset contains a substantial number of “dumps” – chats that were started anew by dumping all the content of a discontinued chat into them. Such chats add redundant information to the dataset, might artificially skew the relation of content posted per author, and distort analyses that leverage the time stamps of messages to reconstruct temporal patterns. To remove such dumps, we therefore excluded chats that either showed activity for a period shorter than 10 days or where more than 90% of the content included was posted in a single day. In total, we excluded 527 such dump chats with 163,056 messages this way. While we do not further use the content from these chats, their sheer existence provides an estimator of how many chats in the domain of conspiracy discussions on Telegram are discontinued. After cleaning the dataset, it contains a total of 56,230,491 messages posted in 6067 chats, covering the time span from September 23, 2015 to August 5, 2022 (see also Tab. 1).
3.0.3 Networks
In addition to the analysis of the content of messages within groups, the Schwurbelarchiv allows for the analysis of the network structure of group affiliation of authors. Viewed through the lens of networks, our dataset can be conceptualised as a network in which groups are nodes and a weighted edge between two nodes indicates how many authors did at least post one message in both groups. In a preliminary analysis, by calculating average degree and edge weight, we find that authors have potential outreach to a mean of 2522.4 other authors (median = 637.0), and share 3.49 common groups with other authors (median = 3.32). Likewise, the data set can be conceptualized as a network in which authors are nodes and a weighted edge between two authors indicates in how many groups both authors participated in the discourse. Here it shows, that a groups shares common at least one common author with a mean average of 1289.0 other groups (median = 1457.0), and groups share a mean of 8.16 authors (median = 2.19). Both networks make it possible to examine topological features of posting activity in the dataset. Based on the messages contained in the Schwurbelarchiv we compiled these networks and make them available for further analysis.
3.0.4 Ethics and Data Protection
While the data contained in the Schwurbelarchiv was collected from publicly accessible chats, it should not be considered “fair game” for research without consideration for the privacy expectations of the users whose messages are contained in the dataset. In preparing the dataset for publication, we considered a number of arguments to weigh the potential infringement on user expectations against the merit of potential research enabled by making the data accessible, which we will detail in this section. Furthermore, we obtained IRB clearance for research with and re-publication of the data contained in the Schwurbelarchiv (vote of the IRB of Graz University of Technology from March 28, 2023).
Telegram users can freely choose their username on the platform. Therefore, more than one user can choose the same username and usernames do not necessarily allow the identification of individuals outside Telegram. Nevertheless, many usernames contained in the meta information of posts resemble names of real people. Even though the content contained in the dataset is public, users of social media platforms usually do not expect their posts to be singled out and discussed in research projects Townsend and Wallace (2016). We therefore argue that any attempt at linking Telegram posts to the identities of people outside the platform would be problematic from the perspective of user expectations. Therefore, we pseudonymised all messages by replacing usernames with unique IDs via hashing. As a result, messages that come from the same username can still be linked, but any resemblance of usernames to real names is removed. We believe that this step does not hinder the exploration of discussion topics and communication dynamics while it helps to balance user expectations with the potential merit of using the Schwurbelarchiv for research.
A similar consideration has to be made for the names of groups and channels in the dataset. While the majority of these names refer to places or discussion topics, a number of chat names resemble the names of real people. This is particularly frequent for channels where single individuals post content that is broadcast to the followers of the channel. For the present data publication, we chose to retain the names of chats and not pseudonymise them. This decision is driven by two main considerations: First, by altering chat names, analyses like the identification of the geographical scope of chats provided in this work would not be possible and our work would not be reproducible. Second, users that choose to create a chat under their name expect to receive a substantial amount of attention. We argue that it is ethically admissible to retain these names Townsend and Wallace (2016), as providing them in the dataset will neither significantly heighten the attention these users receive nor breach their expectations.
Even though usernames are anonymised in our dataset, the content of the posts might allow for the re-identification of people based on disclosed attributes such as location, demographic characteristics or real names that are contained in message texts rather than usernames. Pseudonymisation of this information in message texts let alone multimedia content is not technically feasible. However, re-identification of individuals would require some effort to parse content and link user identities to messages. While we consider this technically feasible for individual messages, we caution against any such attempts to honour expectations of individual users.
While the anonymous collector of the Schwurbelarchiv explicitly states their intention for the usage of the dataset as a source for journalists and research on the website of the dataset [^4] https://web.archive.org/web/20241126233137/
https://schwurbelarchiv.wordpress.com/
, neither the website nor the Internet Archive collection contain a license. This in principle precludes re-use and re-publication of the dataset. However, in private conversation with the authors of the present work, the anonymous collector confirms that they intended to publish their data under a CC0 license and explicitly allow for re-use and re-publication.
4 Results
4.1 Descriptive Statistics
Table 1 shows descriptive statistics of the cleaned Schwurbelarchiv dataset. With the number of channels and groups being almost the same, our dataset can for example be used to investigate whether conspiracy theory related discussion arises organically through conversation in discussion groups, or whether agenda-setting by a small number of users in broadcast channels is more prevalent. We find that media files like images, videos, and voice messages make up 26.4% of all messages. While they account for 38.1% of all forwarded messages, for original messages only 16.9% are media files. This difference makes the analysis of the content of media files particularly relevant for investigating content dissemination. The distinction between forwarded messages and original messages also serves to estimate the scope of the conspiracy related German-language Telegram discourse, which we will leverage in Section “Validation and Completeness” below to gauge the completeness of the dataset.
Original Forwarded Total
Chats 6076
Authors 357,048 96,037 402,549
Channels 3130
Groups 2946
Messages total 56,230,491
Text 36,097,935 12,612,543 48,710,478
Images 5,707,134 5,994,015 11,701,149
Video 1,066,416 1,408,752 2,475,168
Voice messages 543,077 345,803 888,880
Table 1: Descriptive statistics of the content of the Schwurbelarchiv.
4.2 Language, time zone and region
Figure 1: Percentage of German messages in chats.
To evaluate the language distribution within the chats in the dataset, we classify all messages with the FastText package Facebook (2016). Filtering out empty, short (less than 20 characters) and URL-only messages we are left with 52% of all messages. From those, German is by far most prevalent, with 92.9% of the messages followed by English with 1.8%. We also examined the distribution of German message proportions across chats, shown in Figure 1. The vast majority of chats (88.7%) contain at least 50% of messages in German, while in 46.9% of chats it exceeds 90%.
As messages on Telegram are not geo-tagged, we have no direct information about the geolocation of the authors that posted messages contained in the dataset. However, we can use other information to narrow the geographic scope of the observations. In Fig. 2 we show the posting pattern on different weekdays and hours of a day using the UTC+1 time zone (e.g., Central Europe). We observe a consistent minimum in posting frequency during the European night hours with maxima in posting frequency in the early evening hours and on Sundays. This observation indicates that the diurnal rhythm of the majority of authors in the dataset is consistent with Central Europe.
Figure 2: Posting frequency aggregated by hour of the day and day of the week in the UTC+1 time zone.
We also conducted this analysis for the chat names included in the datasets provided by Zehring and Domahidi (2023) and Mohr (2023). Table 2 compares the regional representation across these datasets, showing the total number and percentages of regionally related chats. The Schwurbelarchiv dataset has the highest proportion of regionally related chats, accounting for 25%, compared to 6% and 10% in the other datasets. The high proportion of regional chats in the Schwurbelarchiv dataset highlights its value for the study of localised communication and group dynamics, as opposed to discourse dominated by large national or transnational political actors or influencers.
Number of Regional Related Chats % of full dataset
Schwurbelarchiv 1,514 24.9 %
Mohr 2,275 9.9 %
Zehring 33 5.7 %
Table 2: Comparison of regional Telegram chats across datasets.
4.3 Validation and Completeness of the Dataset
In the light of frequent message deletions on Telegram Buehling (2024) and the absence of a structured database of Telegram content that can be used to retrieve a comprehensive dataset for a given query, it is important to assess the completeness of collected Telegram datasets. To this end, we consider completeness on three levels. On the level of messages, we assess the fraction of deleted content. On the level of forwarded messages, we assess the completeness of the dataset by analysing the presence of original messages and forwarded messages: if many forwarded messages are not present in the dataset as original messages, this is evidence that a substantial part of the discourse is not captured. Lastly, we assess completeness on the level of chats included in the dataset: there is no official record of all German-speaking Telegram chats dedicated to conspiracy-related content. We therefore compare the Schwurbelarchiv with similar datasets, specifically those by Zehring and Domahidi (2023), Mohr (2023) and Baumgartner et al. (2020) introduced earlier.
4.3.1 Deleted Messages
Assessing the completeness of the dataset on a message level involves examining the potential for missing or deleted messages, which is closely tied to the data collection approach. There is no published metadata on the time of data collection. In particular, the start of the streaming process for each chat is not known. However, the difference between bulk download and streaming has direct consequences for the expected completeness of messages in the dataset at different points in time: In the bulk download phase, it is probable that the dataset is missing messages that were deleted by authors prior to the download and that the percentage of missing messages increases further back in time from the time of the download. These deleted messages have no placeholders in the dataset and are therefore not visible. In contrast, during the continual streaming phase, messages are captured in near real-time, therefore reducing the probability of missing messages caused by user deletions. We were not able to detect the transition point where the data collection method changed from bulk to stream download for the different chats. Therefore, we cannot guarantee the completeness of the dataset on a message level, since we cannot indicate the presence or fraction of deleted messages. However, it is probable that message deletions are less likely to have occurred for the later years contained in the dataset.
As stated by Buehling (2024), researchers should consider the systematic bias and possibly skewed results in computational analysis introduced by message deletions when working with such datasets. These findings indicate a key limitation of our dataset, since the lack of metadata detailing the data collection process prevents us from determining the extent to which deleted messages may have influenced its completeness.
4.3.2 Forwarded Messages
To further assess the completeness of conspiracy-related chats in the dataset, we performed a comparative analysis of forwarded messages and messages. In the context of Telegram, a forwarded message refers to a message which has been forwarded and is marked as such in the chat. These forwarded messages often serve as proxies for content influence, reflecting the flow of information between chats. We calculated the proportion of forwarded messages which are also present in the dataset as original messages. This metric indicates the degree to which the original sources of forwarded messages are represented in the dataset. If this value were 100%, it would imply that all forwarded messages are also present in the original chat the message was posted in, achieving complete coverage of the source population. In our dataset, this proportion was calculated to be 47.9%. This finding suggests that we cover about half of the German-language conspiracy related discourse on Telegram in the observation time frame, highlighting the dataset’s limitations in capturing the full extent of the discourse.
4.3.3 Chats
Lastly, we also compare the contents of the Schwurbelarchiv to the chats collected by Zehring and Domahidi (2023), Mohr (2023) and Baumgartner et al. (2020). The dataset by Zehring and Domahidi (2023), consists of 578 chats and covers the time period October 28, 2015, to January 3, 2022. Among these chats, there are two that we have excluded from our dataset due to their classification as dump chats. Out of the remaining chats, 121 are also present in our dataset, representing approximately 22% of Zehring’s chats and 2% of the chats contained in the Schwurbelarchiv. Although this overlap is smaller than expected given the similar selection criteria and seed topics, several factors may account for the discrepancy: Zehring and Domahidi (2023) concentrated their seed selection on groups with substantial societal influence, including opinion leaders, politicians, and well-known conspiracy theorists. In contrast, the Schwurbelarchiv dataset includes numerous smaller, more specialized groups often focused on local issues. For example, groups such as Corona Rebellen have multiple sub-groups tailored to different cities and regions, reflecting a tendency towards decentralization in the Schwurbelarchiv dataset. This is also reflected in the overall share of content we identified as coming from regional groups, which is much higher than for Zehring and Domahidi (2023) (see also Tab. 2).
The subset of German-language chats in the dataset collected by Mohr (2023) which was used for comparison here consists of 22,968 chats. Among these, 46 have been categorised by us as dump chats and have been excluded. Another 801 of the chats in this dataset are also contained in the Schwurbelarchiv. This corresponds to 13% of the chats we have in our dataset, while only comprising 3% of chats contained in the dataset collected by Mohr (2023).
The Pushshift dataset Baumgartner et al. (2020) is the most comprehensive public dataset to date, encompassing over 27,000 chats from various countries. Of these, 78 chats (1.3%) overlap with the Schwurbelarchiv. None of the excluded dump chats in the Schwurbelarchiv are present in the Pushshift dataset. This low overlap is likely due to linguistic and regional differences in the selection criteria: while the Schwurbelarchiv and datasets by Zehring and Domahidi (2023) and Mohr (2023) focus on German-speaking chats, the Pushshift dataset has a more global scope. In addition, the temporal scope of data collection between the Pushshift data set (2015-2019) and the Schwurbelarchiv (2016-2022) differs substantially.
Number of Chats Overlap with Schwurbelarchiv Overlap in %
Mohr (2023) 22,968 801 13 %
Zehring and Domahidi (2023) 578 121 2 %
Baumgartner et al. (2020) 27,801 78 1.3 %
Table 3: Number of chats per dataset and overlap with the 5,963 chats in the Schwurbelarchiv.
Table 3 summarizes the number of chats in each dataset and their overlap with the Schwurbelarchiv, reported both in absolute numbers and as a percentage of Schwurbelarchiv chats. Furthermore, the overlap between the Zehring and Domahidi (2023) and Mohr (2023) datasets is 369 chats, corresponding to 66% and 1.6% of the content of each of these datasets, respectively. A total of 102 chats are present in all three data sets. Additionally, 39 chats are shared between the Mohr (2023) and Pushshift datasets, while no chats overlap between the datasets by Zehring and Domahidi (2023) and Pushshift.
Figure 3: Number of messages (blue) and active authors (red) per day in the Schwurbelarchiv. Maxima in messaging activity and related real-life events were identified manually.
Figure 4: Number of messages in the time period between January 1, 2022 and the end of the observation period (blue) and number of messages posted per person per day (green).
4.4 Overall Posting Activity
Figure 3 shows the overall posting frequency in terms of messages and active chats over the full observation period. Active chats in this context are defined as chats that had at least one posted message on a given day. The number of active chats and message volume already started to increase slightly before the onset of the COVID-19 pandemic. However, the dramatic increase in active users and messages coincides with the onset of the pandemic in Europe. Peaks in message volume coincide with political events such as a large demonstration at the German parliament on August 29, 2022, and the Capitol Riot in the U.S. on January 6, 2021. Further insight into the discussion topics at these times could be gained by employing topic modelling approaches to the messages posted during these events.
In Fig. 4 we again show the overall number of messages per day, overlaid with the number of messages per author and day. After an initial period of high activity in early 2020 with around 13 messages per author and day we observe a steep decline of individual author activity in April and May 2020 that does not coincide with a decline in the overall posting traffic. This indicates that while individual authors became less active, more authors joined the discourse on Telegram, contributing to a steady increase of overall message volume. This is also consistent with our impressions from deep-reading small fractions of the content: we find that as conspiracy-related Telegram chats gain in popularity, established authors are replaced by newcomers that want to learn about conspiracy theories like QAnon, Pizzagate, or Chemtrails, but rarely establish themselves as experts. We call this behaviour Conspiracy Tourism which might be an interesting avenue for further more qualitative research. We observe a more gradual decline in author activity from June 2020 to June 2021 where authors post on average nine messages a day, followed by another steep decline in late summer 2021. At the end of our observation period, e.g. in summer 2022, the activity of authors seems to have stabilised around three messages per author and day.
4.5 Activity distribution
Figure 5: A. Distribution of messages and authors per chat. Full lines indicate the logbinned distributions. B. Cumulative fraction of messages per top chats and authors. $G$ indicates the Gini coefficient of the cumulative distribution.
A common aspect of social media platforms is high inequality, where few drivers (authors, chats) are responsible for most of the content/attention. In Fig. 5 we show this to be true in the Schwurbelarchiv as well. The probability distributions of messages and unique authors per chat (Fig. 5A) have very wide ranges, with the largest chats having on the order of $10^{4}$ authors and $10^{6}$ messages. At the same time, most chats have less than $10$ messages and authors. Due to the uncertainty around the dataset sampling scheme, we refrain from e.g. fitting power-law distributions as the sampling can heavily bias both measured exponents and model selection of the best-fitted distribution (power-law vs lognormal). Regardless, the distributions are clearly heavy-tailed.
Grouping the messages by unique authors and chats (Fig. 5B), we see that $80\%$ of all messages belong to the top $12.1\%$ and $3.1\%$ of chats and authors, respectively. This is reflected in the very high Gini indexes of $G=0.85$ (chats) and $G=0.94$ (authors). This high level of concentration is comparable with what is observed in other social media platforms Orellana-Rodriguez and Keane (2018); Zhu and Lerman (2016).
4.6 Transcription
At the time of this publication, transcription of all files with audio tracks is ongoing. New files will be iteratively added to the online dataset. For the transcription process, we used the openai/whisper-large-v3-turbo model, which was the latest whisper model available at the time of publication Radford et al. (2022). To boost the transcription speed, we combined the model with FlashAttention-2 Dao (2024) which resulted in a transcription rate of around 500x realtime. This increases the amount of text in our dataset from 1,490M words in the written messages to a total of approximately 2,103M words (estimation of video file words, see Table 4) the text corpus by 41%. As media files comprise a considerable share of all forwarded messages, transcribed messages are an important carrier of information between chats. A feature of the text generated via transcription is that it consists of a well-defined vocabulary and therefore has no typos or alternative spellings. While this precludes the analysis of textual characteristics related to these features, it might allow for more accurate text analysis and topic identification and requires less pre-processing before further analysis.
Audio Video Images
Count 0.889M 2.475M 11.71M
Total words 217.8M 395.1M*
Total filesize 15.1TB 20.2TB 1.14TB
Total duration 63,395 h 63,194 h
Median duration 74.96 s 42.57 s
Table 4: Statistics of transcribable media files. Video word count was estimated from a random sample of 6.2K videos.
Next to voice messages and video files, 77% of the 15,065,197 media files that were sent by Telegram users in the dataset were images. Manual investigation of a small fraction of the images suggests that many of them contain text in form of memes, slogans, or announcements. While we believe it technically feasible to extract text from these images or even use new multimodal large language models to create text descriptions of them, such processing incurs additional computational processing cost.
5 Discussion and Conclusions
In this paper, we present the Schwurbelarchiv dataset, a large-scale collection of more than 6,000 Telegram chats collected via snowball-sampling, focusing primarily on German-language discussions in the realm of conspiracy theories, misinformation, and political extremism in the time from 2016 to 2022. Notably, posting activity in the dataset peaks during the COVID-19 pandemic and at key political events of the time, indicating heightened and potentially coordinated online activity in these special interest groups.
With the present data publication, we provide a cleaned and validated version of the dataset that is easy to access and re-use for other researchers. We clarify questions around the ethics of using the dataset as well as the terms of use. The dataset contains digital trace data of individuals. While we pseudonymised usernames, the dataset is not completely anonymised and in principle allows for the re-identification of individuals based on publicly disclosed information contained in the data. We release this dataset because we believe the benefit of being able to study the spread of conspiracy theories in a completely unregulated online environment at scale during a time of crisis outweighs the risk of infringing on user expectations and privacy. However, we do not intend this dataset to be used for the profiling of individuals and strongly caution against any attempts of re-identification.
Rather than relying on data collection from the Telegram API, the dataset was scraped with an automated approach via Telegram desktop clients that enabled the download of multimedia content, e.g. images, videos and voice messages. While the raw dataset was accessible in the Internet Archive before, we make it easier to use for research by providing an assessment of its completeness and scope, parsing and cleaning it. Furthermore and next to the large corpus of text messages, we provide transcriptions of voice messages and videos, enriching the corpus with more than three million files which we make accessible for further analysis. The Schwurbelarchiv is – to our knowledge – the only large-scale Telegram dataset that includes transcribed audio content. Based on our qualitative observations of the chats in the dataset, the discussion is often driven by video or voice message contributions while text messages are mostly used to comment on this multimedia content. Therefore, by ignoring multimedia content – as is often-times the case in research that uses digital trace data – large parts of the discussion are missed. With our work we therefore aim to enable research into the role of multimodal content in spreading conspiratorial narratives.
In terms of information contained in the Schwurbelarchiv, one substantial limitation of our dataset is the lack of unique user identifiers. As users can freely choose their usernames, the lack of access to user IDs makes it impossible to determine with certainty whether different authors are truly distinct individuals or whether posts with the same username originate from the same user. However, we believe the assumption that different usernames largely indicate activity by different individuals is warranted. To address this constraint, we plan to investigate text features and variance in temporal messaging patterns in the future to distinguish between authors who may share the same name.
We assessed the completeness of our dataset by comparing it to the content of comparable Telegram datasets Baumgartner et al. (2020); Mohr (2023); Zehring and Domahidi (2023). This comparison also serves to estimate the scope of conspiracy-related discourse on Telegram during the COVID-19 pandemic in German speaking countries. On the chat level the overlap with other datasets is low, with a maximum of 13% of the chats shared with Mohr (2023). This low overlap can largely be explained by a difference in inclusion criteria used to select chats in the different dataset, and differences in the temporal scope of the data sets. Another way to gauge the completeness of our dataset regarding conspiracy-related discourse is to look at message forwards. Here, we found that 47.9% of incoming forwarded messages are also contained as original messages in our dataset. On the flip side, this means that around half of the forwarded messages originate in chats that are not contained in our dataset. We therefore estimate that we cover around half of the conspiracy-related discourse happening in predominantly German-language Telegram chats at the time in our dataset. As we have no reason to believe that the selection of chats included in our dataset is systematically biased in any way, other than being centred on conspiracy-related discourse in German, we believe that the Schwurbelarchiv is an adequate resource for the study of conspiracy theories in the given temporal and geographic scope.
That being said, another limitation of our dataset relates to the unknown point in time at which data was collected for each chat. As has been reported by Buehling (2024), messages on Telegram are deleted relatively frequently, and deletion patterns are not random. We estimate large portions of our dataset, particularly towards the end of the observation period, to have been collected via streaming and therefore be relatively unaffected by message deletion. However, we cannot determine the point at which streaming started for each chat. Nevertheless, we note that the other datasets covering a similar geographical and temporal scope have all been collected by bulk download through the Telegram API Zehring and Domahidi (2023); Mohr (2023); Baumgartner et al. (2020). Therefore, these data sets are subject to increasing message deletion the farther a point in time is away from the time of data collection as well. By including both bulk-downloaded and streamed portions of chat content, the Schwurbelarchiv is likely more complete in terms of deleted messages than other data sets.
Next to the content of the messages, our data offers insights into the dynamics of information sharing and highlights how message frequency, group activity, and author engagement are distributed and evolve over time. The activity of authors in groups exhibits heavy-tailed distributions that approximate a power-law. This is consistent with activity patterns across social media platforms, such as Twitter Orellana-Rodriguez and Keane (2018). Furthermore, activity distribution is very uneven, with 80% of messages being posted in the top 12.1% of chats 80% of messages being posted by the top 3% of authors.
The Schwurbelarchiv data can also be analysed from a network perspective by treating groups or authors as nodes and linking them according to shared membership or participation. This approach reveals topological features such as common affiliations or frequently co-occurring authors. Future research could harness these networks to investigate community structures, track the spread of (mis-)information and narratives, and identify influential participants across different discussions.
In summary, the Schwurbelarchiv provides broad utility for research in social network analysis, linguistic analysis, political science, and information studies, particularly in the realm of conspiracy theories. Future work can build on the dataset to investigate the dynamics of misinformation spread, conduct topic modelling, explore the influence of real-world events on user behaviour in online communities, and analyse the dynamics of conspiracy-related discourse in individual chats in a more qualitative way.