### v3

# The Schwurbelarchiv: A German-Language Dataset Derived from Multimodal Telegram Content for the Study of Conspiracy Theories

Abstract
Sociality borne by language, as is the predominant digital trace on text-based social media platforms, harbours the raw material for exploring a multitude of social phenomena. Distinctively, the messaging service Telegram provides functionalities that allow for socially interactive as well as one-to-many communication. Our Telegram dataset contains over 5,800 groups and channels and 63 million messages, originating from a data-hoarding initiative named the “Schwurbelarchiv” (from German schwurbeln: speaking nonsense). Uniquely, it includes the transcriptions of over 3 million audio and video files. While the raw data was previously archived on the Internet Archive by an anonymous data hoarder, it was stored in a format that is difficult to process and largely inaccessible for systematic research. Our contribution consists of parsing, cleaning, and validating this raw archive, pseudonymising user data, and transcribing roughly 126,000 hours of audio and video content, thereby transforming this data hoard into a structured, research-ready dataset. This dataset publication details the structure, scope, and methodological specifics of the Schwurbelarchiv, emphasising its relevance for further research on the German-language conspiracy-theory-related discourse. We validate its predominantly German origin by linguistic and temporal markers and situate it within the context of similar datasets. We describe process and extent of the transcription of multimedia files. Thanks to this effort the dataset uniquely supports analysis of text from originally multimodal sources like voice messages and videos to investigate online social dynamics and content dissemination. Researchers can employ this resource to explore societal dynamics related to misinformation, political extremism, opinion adaptation, and social network structures.
keywords:
telegram dataset, conspiracy theories, social media, misinformation
1 Introduction
The proliferation of digital trace data from social media platforms has enriched the study of societal phenomena, enabling researchers to analyse communication patterns and the dissemination of information at large scales [1]. Platforms like Telegram, with its blend of interactive groups and broadcast channels, serve as fertile ground for examining issues ranging from political discourse to the spread of conspiracy theories. The insights gained from such analyses reveal their influence on societal narratives.
The Schwurbelarchiv [^1] The term schwurbeln, while denoting pseudoscientific and irrational content more generally, gained widespread usage in German-speaking countries from 2021 onward in the context of the COVID-19 pandemic. dataset introduced here contributes to a growing number of published Telegram datasets for research, e.g. La Morgia et al. [2], Baumgartner et al. [3], and – with the transcription of audio files – offers a unique starting point to widen the scope of research into conspiracy theories by including multimedia content. This publication aims to provide a resource for researchers to enable large-scale investigations of the societal and linguistic dynamics that are connected with conspiracy theories, misinformation, political dissatisfaction, and extremism online, which has become a growing research interest in recent years [4, 5, 6, 7, 8, 9].
For empirical research on conspiracy theories in German-language in particular, digital trace data from Social Media Platforms or Messenger Services have substantially contributed to the research landscape but still face limitations that we aim to address with the data set described here.  Pustet et al. [10] study conspiracy content in approximately 4,000 German Telegram messages and identify the scarcity of large-scale German data as a limitation for the robustness of the research and its applicability to the broader discourse. They also note that existing datasets are predominantly in English and often keyword-filtered instead of snowball-sampled, which creates systematic biases towards a few particular conspiracy theories and narratives. Next to limitations of scale and sampling, the exclusive reliance on textual data leaves a substantial portion of Telegram communication unstudied. In our dataset, media files such as images, videos, and voice messages make up 24.4% of all messages.  Hartwig et al. [11] explicitly identify the neglect of voice messages in misinformation research as a critical gap, pointing to the absence of accessible voice message databases despite the growing role of audio content in the dissemination of conspiracy theories on platforms like Telegram and WhatsApp. Similarly, Steffen [12] takes a first step towards multimodal analysis of German conspiracy channels by applying topic modelling to text-image pairs, yet notes that audio content remains entirely outside the scope of current approaches, despite its prevalence on Telegram. The Schwurbelarchiv addresses these limitations by providing a large-scale, snowball-sampled, German-language dataset of over 63 million messages from more than 5,800 Telegram chats, combined with the transcription of roughly 126,000 hours of audio and video content, which allows text analyses of multimodal sources that move beyond the text-only, small-scale, and English-centric resources currently available.
We emphasise that the raw data contained in the Schwurbelarchiv was not collected by the authors of this paper but was assembled by an anonymous individual and deposited on the Internet Archive. In its original form, the archive comprises nearly 24 TB of compressed multimedia files in formats that are not readily accessible for computational analysis. The contribution of the our work is therefore not the data collection itself, but the substantial effort of making this data usable for research: we parse and clean the raw archive, pseudonymise user information, assess the dataset’s coverage, and transcribe over three million audio and video files. The result is a structured, text-based dataset that can be directly used for computational social science research.
1.1 Telegram.
Telegram is an instant messaging platform where users can exchange messages by creating an account with their phone number. Users can communicate in direct chats (one-to-one communication), in groups of up to 200,000 users (many-to-many), or participate in broadcast channels (one-to-many). These broadcast channels allow an unlimited amount of Telegram users to follow a single Telegram user’s (the administrator’s) content, and therefore broadcast channels are well designed for information dissemination at a large scale. While one-to-one communication is private and generally not accessible to researchers, groups and channels are public and accessible if their URL is known. We refer to groups and channels together as “chats” in the remainder of this work. Notably, Telegram does not offer a global search function for discovering chats by topic, which necessitates snowball sampling approaches as employed by the individual that curated the dataset described here.
2 Related Works
A number of studies on Telegram datasets or research primarily using Telegram data have been conducted in recent years, many of which focus on the dissemination of misinformation and conspiracy theories in online environments. Here, we introduce the three Telegram datasets most closely related to our work in terms of temporal and linguistic scope of the data contained. We analyse similarities and differences across datasets, aiming to provide an assessment of the coverage of the Schwurbelarchiv and situating the dataset within the relevant research context.
2.1 Public Telegram Datasets.
The study by  Zehring and Domahidi [9] investigates the connections between the Querdenken movement, a prominent anti-COVID protest network in German-speaking countries often linked to conspiracy theories, to far-right and alternative media, focusing on content-sharing patterns and the influence of QAnon within the network. Using a dataset of 6,294,955 text messages from 578 German public Telegram channels, the research explores the movement’s key topics over time and across its sub-communities. The dataset was collected by first identifying seed channels by searching for actors connected to the Querdenken and far-right movements, to the esoteric/spirituality scene as well as nominees for the Golden Tin Foil Hat prize  [^2] https://www.dergoldenealuhut.de in Germany, Austria, and Switzerland. This approach resulted in 238 initial channels, from which all messages were scraped. Next, frequently forwarded channels (mentioned 50 times or more) were identified and included in the dataset. This procedure – also called “snowball sampling” – was performed twice and led to a total of 578 public Telegram channels from which all messages available for the time between October 28, 2015 and January 3, 2022 were collected. By performing topic modelling on the dataset, Zehring and Domahidi [9] found four overarching topics present in the Telegram channels: promotion, QAnon, right-wing populism, and COVID-19 conspiracy theories.
A larger, at the time of writing unpublished, dataset on similar topics was assembled by Mohr [13]. Since snowball sampling on its own quickly yields an unmanageable amount of chats, they implemented an automated guidance mechanism that prioritises chats likely to contain relevant data based on a keyword list translated into 44 languages. The keywords targeted content containing misinformation related to COVID-19. The dataset, which is to our knowledge of the largest Telegram datasets available, includes 433,000 chats, comprising 2.1 billion messages, primarily in Russian, English, and German. The dataset contains text, metadata, and links, but excludes videos, images, and voice messages, which is a limitation that is central to the distinction with the Schwurbelarchiv. The messages predominantly reflect activity from European time zones in the period from January 2020 to January 2023, with trends linked to events such as the European COVID-19 vaccination campaign and the start of the Russian invasion of Ukraine. All predominantly German-language chats (23,000) from this dataset were made available to the authors of this study for comparison to the Schwurbelarchiv.
The dataset collected by Baumgartner et al. [3], the largest publicly available collection of Telegram data to date, includes 27,000 chats and 317 million messages. The dataset is available for the years 2015-2019 and was regularly updated during that time, with source code available to facilitate further data collection, enabling researchers to extract information from specific chats. Starting with a seed list of 250 English-language chats – 124 focused on right-wing extremism and 137 on cryptocurrency – the number of chats included was also expanded using a snowball sampling approach. As for the dataset by Mohr [13], this data set includes textual data only. Table 1 provides a summary comparison of these datasets.
Schwurbelarchiv Zehring & Domahidi [9] Mohr [13] Pushshift [3] TGDataset [2]
Chats 5,807 578 22,968 27,801 120,979
Messages 63M 6.3M 2.1B∗∗ 317M 498M
Time span 2015–2022 2015–2022 2020–2023 2015–2019 2015–2022
Language German German Multilingual Multilingual Multilingual
Topic focus Conspiracy Conspiracy / far-right COVID misinfo General General
Transcriptions Yes (126,839 h) No No No No
Publicly available Upon request No No Yes Yes
Table 1: Comparison of Telegram datasets. German-language subset; the full dataset contains 128,000 chats. ∗∗Total across all languages. Channels only; no groups.
2.2 Research on German-Language Conspiracy Content on Telegram.
Beyond the collection of datasets, a number of studies have used German-language Telegram data to investigate conspiracy-related content using digital-trace data and computational methods.  Weigand et al. [14] performed a long-term content analysis of conspiracy narratives in the German Querdenken protest movement, fine-tuning a DistilBERT model on self-annotated data from 34 chat groups comprising 822,000 text messages.  Pustet et al. [10] compared supervised fine-tuning of BERT-based models with prompt-based approaches using GPT-4 and Llama 2 for detecting conspiracy theories in German Telegram messages, building on an annotated corpus of approximately 4,000 messages. More recently,  Steffen [12] applied topic modelling on text-image pairs from 571 German conspiracy channels but still ignores voice messages and videos. These studies share a common set of limitations that our dataset helps to overcome: They operate on comparatively small corpora and are limited to text or at most text and images. Audio and video content constitute a considerable share of Telegram messages, but has not been incorporated into any of these analyses. The Schwurbelarchiv, with its scale and the transcription of audio files, is designed to overcome several of these limitations. For instance, the dataset enables investigating how conspiracy narratives evolve across thousands of groups over time, examining the role of voice messages and videos in disseminating conspiracy content compared to text, or tracing information flows through forwarding networks to understand how specific narratives spread between groups.
3 Methodology
3.1 Data Collection.
The entire raw data contained in the Schwurbelarchiv is available on the Internet Archive [^3] https://archive.org/details/schwurbel-archiv [^4] https://web.archive.org/web/20241126233137/https://schwurbelarchiv.wordpress.com/. The person initially responsible for collecting the data remains anonymous, but responded to requests via email by the authors and provided information about the data selection process. They describe their motivation for collecting the dataset as the pure joy of data-hoarding, contribution to the Internet Archive, and enablement of research and journalism using the collected data as a source.
Similar to the other Telegram datasets described here, groups and channels in the Schwurbelarchiv were selected via a snowball sampling approach. Starting with a seed set of Telegram chats focused on renowned German-speaking conspiracy theorists, other chats were identified via invitation links and forwarded messages originating from the chats in the sample. Newly discovered chats were added to the sample, and all messages posted in them up to the point of their discovery were scraped via bulk download, after which a continual streaming process captured new content in near real-time. Relevance of newly discovered chats was determined by manual curation (e.g., human in the loop) that excluded chats deemed unrelated to conspiracy theories in order to prune irrelevant sampling paths early. Data collection was terminated when the anonymous collector lacked the resources to continue. The resulting dataset covers the time span from September 23, 2015 to August 5, 2022.
Snowball sampling promises to capture a more representative portion of the discourse but still has a number of limitations: manual curation of discovered chats ensures relevance but remains subjective, relevant chats might not be included because sampling terminated before a link to them was found, and relevant chats might be missed because they are entirely disconnected from the forwarding network in which sampling was initiated. In particular, chats from which few messages are forwarded are systematically less likely to be discovered by this approach. We note, however, that snowball sampling is the only available methodology for data collection from Telegram beyond manual search and selection, as the platform lacks a searchable index of public chats.
In contrast to the data collection approach of the other studies described here which relied on the Telegram API, the data in the Schwurbelarchiv was gathered using a Windows Virtual Machine, with four separate Telegram Desktop sessions, each operating under its own account. Chats were then exported using the Remote Desktop Protocol. Once a chat has been discovered by the snowball sampling process, all it’s content up to that point in time was bulk-downloaded. Notably, this approach also enables streaming of the content after a chat has been discovered. This is important because the deletion of messages by users on Telegram has emerged as a significant factor influencing the completeness and reliability of Telegram datasets. Buehling [15] highlight how this can lead to systematic biases in analyses such as topic modelling, dictionary-based computational content analysis, and network studies. The study finds that for channels, on average, only 88% of messages remain accessible five days after posting, dropping to 83% after seven months. In groups, the decline is even steeper, with only 64% of messages remaining after five days and 52% after seven months, emphasizing the need for timely data collection.
The Rempote Desktop Protocol enables storing not only the HTML text files, but all video files, voice messages, and images posted in the chats. Video files larger than 8 MB have not been downloaded until mid of 2021, whereafter the limit for downloaded videos was increased to 50 MB. To our knowledge, the Schwurbelarchiv is the only publicly available Telegram dataset that contains multimedia content. This content constitutes a large part of the discourse taking place on social media platforms that is typically ignored due to the difficulty of handling diverse data formats and large data volumes. The fact that our dataset retains this content and makes the multimodal files accessible through transcribing them into text-data greatly expands the types of research questions and analysis approaches that it enables.
All data columns describing message content and metadata (see Table 5) were parsed from the HTML files exported by the original data collector. The columns message_lang, audio_lang, and video_lang were added by us using the Lingua language detection package  [^5] https://github.com/pemistahl/lingua-py, is_conspiracy was computed using the fine-tuned BERT model by Pustet et al. [10] as described in Section “Conspiracy Content in Telegram Chats” below, and all transcription-related columns (audio_transcription, video_transcription, and associated duration and file size fields) were generated by us using the Whisper model described in the following section.
3.2 Data Cleaning and Preprocessing
Since the raw collected data contains multimedia files, it requires a storage space of nearly 24 TB in a compressed state. The sheer size of the data as well as the native Internet Archive file format .warc complicate processing and analysis. Therefore, next to clarifying ethical, data protection-related, and licensing questions associated with the use of the Schwurbelarchiv for research (see Section “Ethics and Data Protection” below), we set out to preprocess the dataset to make it more easily accessible for a wider research community. This preprocessing involved data cleaning, transcription of audio content, and an assessment of the completeness of the dataset. We describe these steps in more detail in the remainder of this section. The fully documented pipeline is available on GitHub [^6] https://github.com/cs2-lab/schwurbelarchiv.
After downloading the raw data from the Internet Archive, we used Python scripts to extract all relevant information, namely author, posting date, and message body. In addition, we distinguish between original messages and forwarded messages – and their respective authors and posting dates – which enables the reconstruction of message forwarding networks within the dataset. Furthermore, we isolate links to websites shared in the message text body. A full listing of all data columns contained in the cleaned dataset is provided in the Appendix 5. While the raw dataset as it is published on the Internet Archive contains uncensored usernames, we pseudonymize usernames using a SHA-256 hash-function. We extract usernames from the HTML files containing the content of the chats but do not have access to user IDs assigned by Telegram. As a result, we cannot distinguish between users that have chosen the same username. This is a substantial limitation of our dataset. We further note that Telegram allows the use of bots as automated accounts, which may also be present in the dataset. We did not attempt to identify or flag bot accounts, which constitutes another limitation of the dataset.
In addition to textual data, we transcribed the audio of both audio and video files. We used ffmpeg to first transcode the various audio codecs into WAV, and openai/whisper-large-v3-turbo [^7] https://huggingface.co/openai/whisper-large-v3-turbo to transcribe the audio content. Various optimizations were used in order to make the transcription task more computationally efficient, including using the CTranslate2 faster-whisper implementation [^8] https://github.com/SYSTRAN/faster-whisper with Voice Activity Detection and batching. These resulted in a transcription real time factor of $\sim 250$X per Nvidia L40S GPU. Overall, 126,839 hours of content (14.9 years) were transcribed.
During the initial data collection, there was no quality control of the included chats with respect to research interests such as the investigation of temporal conversation patterns. However, the raw dataset contains a substantial number of “dumps” – chats that were started anew by dumping all the content of a discontinued chat into them. Such chats add redundant information to the dataset, might artificially skew the relation of content posted per author, and distort analyses that leverage the time stamps of messages to reconstruct temporal patterns. To remove such dumps, we therefore excluded chats that either showed activity for a period shorter than 10 days or where more than 90% of the content included was posted in a single day. In total, we excluded 1,050 such dump chats (13.7% of total) with 317,532 messages (0.5% of total). While we do not further use the content from these chats, their sheer existence provides an estimation of how many chats in the domain of conspiracy discussions on Telegram are discontinued. After cleaning the dataset, it contains a total of 63,908,365 messages posted in 5807 unique chats, covering the time span from 23/09/2015 to 05/08/2022 (see also Tab. 2).
3.3 Ethics and Data Protection
While the data contained in the Schwurbelarchiv was collected from publicly accessible chats, it should not be considered “fair game” for research without consideration for the privacy expectations of the users whose messages are contained in the dataset. In preparing the dataset for publication, we considered a number of arguments to weigh the potential infringement on user expectations against the merit of potential research enabled by making the data accessible, which we will detail in this section.
Telegram users can freely choose their username on the platform. Therefore, more than one user can choose the same username and usernames do not necessarily allow the identification of individuals outside Telegram. Nevertheless, many usernames contained in the meta information of posts resemble names of real people. Even though the content contained in the dataset is public, users of social media platforms usually do not expect their posts to be singled out and discussed in research projects [16]. We therefore argue that any attempt at linking Telegram posts to the identities of people outside the platform would be problematic from the perspective of user expectations. Therefore, we pseudonymized all messages by replacing usernames with unique IDs via hashing. As a result, messages that come from the same username can still be linked, but any resemblance of usernames to real names is removed. We believe that this step does not hinder the exploration of discussion topics and communication dynamics while it helps to balance user expectations with the potential merit of using the Schwurbelarchiv for research.
A similar consideration has to be made for the names of groups and channels in the dataset. While the majority of these names refer to places or discussion topics, a number of chat names resemble the names of real people. This is particularly frequent for channels where single individuals post content that is broadcast to the followers of the channel. For the present data publication, we chose to retain the names of chats and not pseudonymize them. This decision is driven by two main considerations: First, by altering chat names, analyses like the identification of the geographical scope of chats provided in this work would not be possible and our work would not be reproducible. Second, users that choose to create a chat under their name expect to receive a substantial amount of attention. We argue that it is ethically admissible to retain these names [16], as providing them in the dataset will neither significantly heighten the attention these users receive nor breach their expectations.
Even though usernames are anonymized in our dataset, the content of the posts might allow for the re-identification of people based on disclosed attributes such as location, demographic characteristics or real names that are contained in message texts rather than usernames. Pseudonymization of this information in message texts let alone multimedia content is not technically feasible. However, re-identification of individuals would require some effort to parse content and link user identities to messages. While we consider this technically feasible for individual messages, we caution against any such attempts to honour expectations of individual users.
While the anonymous collector of the Schwurbelarchiv explicitly states their intention for the usage of the dataset as a source for journalists and research on the website of the dataset [^9] https://web.archive.org/web/20241126233137/, neither the website nor the Internet Archive collection contain a license. This in principle precludes re-use and re-publication of the dataset. However, in private conversation with the authors of the present work, the anonymous collector confirms that they intended to publish their data under a CC0 license and explicitly allow for re-use and re-publication.
Furthermore, we obtained IRB clearance for research with and re-publication of the data contained in the Schwurbelarchiv (vote of the Institutional Review Board of Graz University of Technology from March 28, 2023).
4 Results
In the following section, we will first describe the Schwurbelarchiv in more detail, giving statistics on the different content types it contains as well as linguistic and temporal markers. In addition, we provide an analysis of overall posting activity and activity distribution across authors and chats. After establishing basic characteristics of the data set, we describe our efforts to assess the completeness of the data set and the extent to which it relates to conspiracy theories. Lastly, we provide more detail on the results of the audio content transcription.
4.1 Descriptive Statistics
Table 2 shows descriptive statistics of the cleaned Schwurbelarchiv dataset. With the ratio of channels to groups in our dataset being roughly 2:1  [^10] We use HTML-parsing heuristics to determine the category of a chat, which resulted in a small number of chats (37) to be manually determined., our dataset can for example be used to investigate whether conspiracy theory related discussion arises organically through conversation in discussion groups, or whether agenda-setting by a small number of users in broadcast channels is more prevalent. We find that media files like images, videos, and voice messages make up 24.4% of all messages. While they account for 42.2% of all forwarded messages, for original messages only 16.7% are media files. This difference makes the analysis of the content of media files particularly relevant for investigating content dissemination. The distinction between forwarded messages and original messages also serves to estimate the scope of the conspiracy related German-language Telegram discourse, which we will leverage in Section “Validation and Completeness” below to gauge the completeness of the dataset.
Original Forwarded Total
Total chats 5,807
Channels 3,486
Groups 2,321
Authors 348,812 611,164 896,339
Content 43,739,502 18,855,979 62,595,481
Text 38,021,236 15,797,326 53,818,562
Images 5,707,321 6,166,370 11,873,691
Video 1,065,146 1,431,740 2,496,886
Voice messages 542,928 354,531 897,459
External links 6,965,776 5,356,078 12,321,854
Table 2: Descriptive statistics of the content of the Schwurbelarchiv. A small number of chats (15) were of an unknown category.
4.2 Language, time zone and region
To evaluate the language distribution within the chats in the dataset, we classify all messages with the Lingua package. Filtering out empty, short (less than 5 characters, e.g. emoji replies) and URL-only messages we are left with 52.5% of all messages. From those, German is by far most prevalent, with 91.3% of the messages followed by English with 2.2%. We also examined the distribution of German message proportions across chats, shown in Figure 1A. The vast majority of chats (98.6%) contain at least 50% of messages in German, while in 85.5% of chats it exceeds 90%.
As messages on Telegram are not geo-tagged, we have no direct information about the geolocation of the authors that posted messages contained in the dataset. However, we can use other information to narrow the geographic scope of the observations. In Fig. 1B we show the posting pattern on different weekdays and hours of a day using the UTC+1 time zone (e.g., Central Europe). We observe a consistent minimum in posting frequency during the European night hours with maxima in posting frequency in the early evening hours and on Sundays. This observation indicates that the diurnal rhythm of the majority of authors in the dataset is consistent with Central Europe.
Figure 1: A. Percentage of German messages in chats. B. Posting frequency aggregated by hour of the day and day of the week in the UTC+1 time zone. C. Frequency distribution of messages and authors per chat. G indicates the Gini coefficient of the cumulative distribution. D. Lorenz plot showing cumulative fraction of messages per top chats and authors.
4.3 Overall Posting Activity
Figure 2A shows the overall posting frequency in terms of messages and active chats over the full observation period. Active chats in this context are defined as chats that had at least one posted message on a given day. The number of active chats and message volume already started to increase slightly before the onset of the COVID-19 pandemic. However, the dramatic increase in active chats and messages coincides with the onset of the pandemic in Europe. Peaks in message volume coincide temporally with political events such as a large demonstration at the German parliament on August 29, 2020, and the Capitol Riot in the U.S. on January 6, 2021. However, these associations are based on temporal co-occurrence and have not been verified through topic modelling or similar content analyses. Further insight into the discussion topics at these times could be gained by employing topic modelling approaches to the messages posted during these events.
In Fig. 2B we again show the overall number of messages per day, overlaid with the number of messages per author and day. After an initial period of high activity in early 2020 with around 13 messages per author and day we observe a steep decline of individual author activity in April and May 2020 that does not coincide with a decline in the overall posting traffic. This indicates that while individual authors became less active, more authors joined the discourse on Telegram, contributing to a steady increase of overall message volume. This is also consistent with our impressions from deep-reading small fractions of the content: we find that as conspiracy-related Telegram chats gain in popularity, established authors are replaced by newcomers that want to learn about conspiracy theories like QAnon, Pizzagate, or Chemtrails, but rarely establish themselves as experts. We call this behaviour Conspiracy Tourism which might be an interesting avenue for further, more qualitative research. We observe a more gradual decline in author activity from June 2020 to June 2021 where authors post on average nine messages a day, followed by another steep decline in late summer 2021. At the end of our observation period, e.g. in summer 2022, the activity of authors seems to have stabilized around three messages per author and day.
Figure 2: A. Number of messages (blue) and active chats (red) per day in the Schwurbelarchiv. Maxima in messaging activity and related real-life events were identified manually. B. Number of messages in the time period between January 1, 2022 and the end of the observation period (blue) and number of messages posted per person per day (green). The numbers in parentheses indicate the position of the events highlighted in panel A.
4.4 Activity distribution
A common aspect of social media platforms is high activity concentration, where few drivers (authors, chats) are responsible for most of the content/attention [17]. The Schwurbelarchiv is no different. Grouping the messages by unique authors and chats (Fig. 1C), we see that the top 10% of chats aggregate 73.3% of all messages and the top 10% authors contribute 94.6% of all messages. This is reflected in the very high Gini indexes of G=0.83 (chats) and G=0.95 (authors). For comparison, Gini coefficients reported for attention metrics on Twitter are in a similar range with G=0.94 for followers and G=0.90 for retweets [18]. It is, however, important to mention that these measure different quantities than posting activity per chat and are therefore not directly comparable. For Telegram and other messaging platforms, comparable activity concentration metrics are, to our knowledge, not available in the published literature.
The probability distributions of messages and unique authors per chat (Fig. 1D) have very wide ranges, with the largest chats having on the order of $10^{4}$ authors and $10^{6}$ messages. Due to the uncertainty around the dataset sampling scheme, we refrain from e.g. fitting power-law distributions as the sampling can heavily bias both measured exponents and model selection of the best-fitted distribution (e.g. power-law vs. lognormal). Regardless, the distributions demonstrate fat or heavy tails.
4.5 Validation and Completeness of the Dataset
In light of the absence of a structured database of Telegram content that can be used to retrieve a comprehensive dataset for a given query, it is important to assess the completeness of collected Telegram datasets and its relatedness to the targeted scope, e.g. German language conspiracy related discourse. Since Telegram does not provide a searchable index of public chats nor any official record of chats about specific topics, no ground truth exists with which the completeness of any Telegram dataset could be measured. All existing Telegram datasets therefore rely on manual search and selection or snowball sampling. As discussed in Section “Data Collection” above, snowball sampling is systematically biased toward including chats with higher forwarding activity, meaning that smaller or more insular chats are likely underrepresented in the dataset. The metrics we report can therefore provide only an approximate assessment of completeness. To this end, we consider alternative means of studying completeness, using the information that is available about the data set and in the published literature: The anonymous collector of the data set states that the Schwurbelarchiv contains conspiracy-related content. We start by validating this claim via classification of message content. Next, and motivated by the reported frequent message deletions on Telegram [15], we gauge the probability of content deletions to have taken place before data was collected given the data collection approach employed. Next, we attempt to gauge the extent to which the data set covers the discourse that is in scope of the sampling. While no searchable index of existing content exists, forwarded content provides a means of assessing the proportion of the targeted discourse that is included in the sample by analysing the presence of original messages and forwarded messages: if many forwarded messages are not present in the dataset as original messages, this is evidence that a substantial part of the discourse is not captured. Lastly, we employ a complementary approach to assess completeness of the data set by looking at the chats that the Schwurbelarchiv includes. Since there is no official record of all German-speaking Telegram chats dedicated to conspiracy-related content we compare the chats in the dataset with chats in similar datasets, specifically those by Zehring and Domahidi [9], Mohr [13] and Baumgartner et al. [3] introduced earlier.
4.5.1 Conspiracy Content in Telegram Chats
As described in the Methods Section above, the dataset was collected anonymously through snowball sampling of conspiracy-related Telegram groups with the primary aim of capturing conspiracy-related content on Telegram. However, other than manual curation by the data collector, no validation was conducted to ensure that the dataset effectively captured conspiracy content, nor was there an assessment of the proportion of messages within the chats that actually contain such content.
To validate the scope of the dataset, we applied a fine-tuned BERT model developed by Pustet et al. [10], which is designed to detect conspiracy-related content in German Telegram messages. This model was applied to all messages exceeding 5 characters in length. The results showed that 18.4% of chats (1055 chats) do not contain conspiracy-related content. However, these chats account for only 0.5% of all messages longer than 150 characters. This indicates that the majority of the dataset indeed contains conspiracy-related material.
4.5.2 Deleted messages
Assessing the completeness of the dataset on a message level involves examining the potential for missing or deleted messages, which is closely tied to the data collection approach. There is no published metadata on the time of data collection. In particular, the start of the streaming process (after the initial bulk download upon discovery) for each chat is not known. However, the difference between bulk download and streaming has direct consequences for the expected completeness of messages in the dataset at different points in time: In the bulk download phase, it is probable that the dataset is missing messages that were deleted by authors prior to the download and that the percentage of missing messages increases further back in time from the time of the download. These deleted messages have no placeholders in the dataset and are therefore not visible. In contrast, during the continual streaming phase, messages are captured in near real-time, therefore reducing the probability of missing messages caused by user deletions. We were not able to detect the transition point where the data collection method changed from bulk to stream download for each chat. Therefore, we cannot guarantee the completeness of the dataset on a message level with respect to message deletions, since we cannot indicate the presence or fraction of deleted messages. However, it is probable that message deletions are less likely to have occurred for the later years contained in the dataset and that the Schwurbelarchiv is less affected by message deletions than other comparable data sets.
Even though our data set likely has less deleted messages, researchers using the data should still consider the systematic bias and possibly skewed results in computational analysis introduced by message deletions, as stated by Buehling [15]. These findings indicate a key limitation of our dataset, since the lack of metadata detailing the data collection process prevents us from determining the extent to which deleted messages may have influenced its completeness.
4.5.3 Forwarded Messages
To further assess the completeness of conspiracy-related chats in the dataset, we performed a comparative analysis of forwarded and original messages. In the context of Telegram, a forwarded message refers to a message which has been forwarded and is marked as such in the chat. These forwarded messages often serve as proxies for content influence, reflecting the flow of information between chats. We calculated the proportion of forwarded messages which are also present in the dataset as original messages. This metric indicates the degree to which the original sources of forwarded messages are represented in the dataset. If this value were 100%, it would imply that all forwarded messages are also present in the original chat the message was posted in, achieving complete coverage of the source population that is in scope of the sampling. In our dataset, this proportion was calculated as the fraction of forwarded messages whose original sender (fwd_author) and original posting time (fwd_posting_date in UTC) jointly match an original (non-forwarded) message in the dataset (matching on author and posting_date after UTC normalisation). In the dataset this value is 54.1%. This finding suggests that we cover about half of the German-language conspiracy related discourse on Telegram in the observation time frame, highlighting the dataset’s limitations in capturing the full extent of the discourse. We interpret this proportion as a rough estimate of coverage. A higher or lower value is not inherently desirable, as it reflects the scope of the sampled forwarding network rather than the quality of the data contained.
4.5.4 Chats Covered
Lastly, we also compare the contents of the Schwurbelarchiv to the chats collected by Zehring and Domahidi [9], Mohr [13] and Baumgartner et al. [3]. The dataset by Zehring and Domahidi [9], consists of 578 chats and covers the time period October 28, 2015, to January 3, 2022. Among these chats, there are two that we have excluded from our dataset due to their classification as dump chats. Out of the remaining chats, 121 are also present in our dataset, representing approximately 22% of Zehring’s chats and 2% of the chats contained in the Schwurbelarchiv. Although this overlap is smaller than expected given the similar selection criteria and seed topics, several factors may account for the discrepancy: Zehring and Domahidi [9] concentrated their seed selection on groups with substantial societal influence, including opinion leaders, politicians, and well-known conspiracy theorists. In contrast, the Schwurbelarchiv dataset includes numerous smaller, more specialized groups often focused on local issues. For example, groups such as Corona Rebellen have multiple sub-groups tailored to different cities and regions, reflecting a tendency towards decentralization in the Schwurbelarchiv dataset.
The subset of German-language chats in the dataset collected by Mohr [13] which was used for comparison here consists of 22,968 chats. Among these, 46 have been categorised by us as dump chats and have been excluded. Another 801 of the chats in this dataset are also contained in the Schwurbelarchiv. This corresponds to 13% of the chats we have in our dataset, while only comprising 3% of chats contained in the dataset collected by Mohr [13].
The Pushshift dataset [3] is the most comprehensive public dataset to date, encompassing over 27,000 chats from various countries. Of these, 78 chats (1.3%) overlap with the Schwurbelarchiv. None of the excluded dump chats in the Schwurbelarchiv are present in the Pushshift dataset. This low overlap is likely due to linguistic and regional differences in the selection criteria: while the Schwurbelarchiv and datasets by Zehring and Domahidi [9] and Mohr [13] focus on German-speaking chats, the Pushshift dataset has a more global scope. In addition, the temporal scope of data collection between the Pushshift data set (2015-2019) and the Schwurbelarchiv (2016-2022) differs substantially.
Table 3 summarizes the number of chats in each dataset and their overlap with the Schwurbelarchiv, reported both in absolute numbers and as a percentage of Schwurbelarchiv chats. Furthermore, the overlap between the Zehring and Domahidi [9] and Mohr [13] datasets is 369 chats, corresponding to 66% and 1.6% of the content of each of these datasets, respectively. A total of 102 chats are present in all three data sets. Additionally, 39 chats are shared between the Mohr [13] and Pushshift datasets, while no chats overlap between the datasets by Zehring and Domahidi [9] and Pushshift.
Dataset Description Number of Chats Overlap Overlap in %
Mohr [13] COVID-19 misinformation; multilingual; keyword-guided snowball sampling 22,968 801 13.0%
Zehring & Domahidi [9] German Querdenken movement; far-right and conspiracy content 578 121 2.0%
Pushshift [3] General; multilingual; seeded from right-wing extremism and cryptocurrency channels 27,801 78 1.3%
Table 3: Number of chats per dataset and overlap with the 5,963 chats in the Schwurbelarchiv.
Additionally, we compared our dataset with the channels and groups monitored by the Machine Against the Rage initiative by BAG Gegen Hass im Netz [19], which shows network patterns of forwarded messages within the ideological spectrum of conspiracy theories and extremism. We found an overlap of 212 chats with the Schwurbelarchiv. The initiative’s monitoring begins right after our observation period in September 2022 and therefore allows for contextualization of these groups’ further embeddedness in the German speaking Telegram network.
4.6 Transcription
Uniquely, the dataset contains transcriptions of both audio messages and video files. For the transcription process, we used the openai/whisper-large-v3-turbo model, which was the latest whisper model available at the time of publication [20]. This increases the amount of text in our dataset from 1,761M words in the written messages to a total of approximately 2,718M words (see Table 4), increasing the text corpus by 54.3%. As media files comprise a considerable share of all forwarded messages, transcribed messages are an important carrier of information between chats. A feature of the text generated via transcription is that it consists of a well-defined vocabulary and therefore has no typos or alternative spellings. While this precludes the analysis of textual characteristics related to these features, it might allow for more accurate text analysis and topic identification and requires less pre-processing before further analysis. Regarding transcription quality, the Whisper large-v3-turbo model used for transcription achieves Word Error Rates (WER) between approximately 4% and 7% for German on standard benchmarks such as Common Voice 15 and Fleurs [20]. However, these benchmarks consist of clean, read-aloud speech and therefore represent a best-case transcription accuracy. The audio content in the Schwurbelarchiv includes potentially incomprehensible speech features like informality, videos with background noise, dialect speech, all of which are likely to increase error rates. To accommodate these challenges, we employed Voice Activity Detection (VAD) to filter non-speech segments, to help mitigate noise from non-voice signals. As for potential improvements: speaker diarization and the use of the full (non-turbo) Whisper model would have offered potential quality improvements but were out of scope due to the substantially higher computational cost. We consider the chosen pipeline to represent the best trade-off between transcription quality and feasibility at this scale. We did not conduct a manual evaluation of transcription quality on a sample of our data, which constitutes a limitation of this work.
Text Voice messages Video Images
Count 54.10M 0.87M 1.58M 11.87M
Total words 1,761M 530M 427M
Total filesize 0.91 TB 17.43 TB 1.14 TB
Total duration 65,433 h 61,406 h
Median duration 76.83 s 74.00 s
Table 4: Multimodal data statistics. Excludes video and audio files that were corrupted, didn’t include an audio track or had an unknown codec.
Next to voice messages and video files, 78% of the 15.3M media files that were sent by Telegram users in the dataset were images. Manual investigation of a small fraction of the images suggests that many of them contain text in form of memes, slogans, or announcements. While we believe it technically feasible to extract text from these images or even use new multimodal large language models to create text descriptions of them, such processing incurs additional computational processing cost and is left to future work.
5 Discussion and Conclusions
In this paper, we present the Schwurbelarchiv dataset, a large-scale collection of more than 5,800 Telegram chats, focusing primarily on German-language discussions in the realm of conspiracy theories, misinformation, and political extremism in the time from 2016 to 2022. Notably, posting activity in the dataset peaks during the COVID-19 pandemic and at key political events of the time, indicating heightened and potentially coordinated online activity in these special interest groups.
With the present data publication, we provide a cleaned and validated version of the dataset that is easy to access and re-use for other researchers. We clarify questions around the ethics of using the dataset as well as the terms of use. The dataset contains digital trace data of individuals. While we pseudonymized usernames, the dataset is not completely anonymised and in principle allows for the re-identification of individuals based on publicly disclosed information contained in the data. We release this dataset because we believe the benefit of being able to study the spread of conspiracy theories in a completely unregulated online environment at scale during a time of crisis outweighs the risk of infringing on user expectations and privacy. In particular, what we add with this publication is data cleaning and validation of a data set that was originally collected by an anonymous data hoarder, which makes the data set more usable for researchers. Malicious use that attempts to single out individuals is already possible with the original data set available on the Internet Archive. However, we emphasize that we do not intend this dataset to be used for the profiling of individuals and strongly caution against any attempts of re-identification.
Rather than relying on data collection from the Telegram API, the dataset was scraped with an automated approach via Telegram desktop clients that enabled the download of multimedia content, e.g. images, videos and voice messages. While the raw dataset was accessible on the Internet Archive before, we make it easier to use for research by providing an assessment of its completeness and scope, parsing and cleaning it. Furthermore and next to the large corpus of text messages, we provide transcriptions of voice messages and videos, enriching the corpus with more than 126,000 hours of transcribed audio and video content which we make accessible for further analysis. The Schwurbelarchiv is – to our knowledge – the only large-scale Telegram dataset that includes transcribed audio content. Based on our qualitative observations of the chats in the dataset, the discussion is often driven by video or voice message contributions while text messages are mostly used to comment on this multimedia content. Therefore, by ignoring multimedia content – as is often-times the case in research that uses digital trace data – large parts of the discussion are missed. With our work we therefore aim to enable research into the role of such content in spreading conspiratorial narratives.
In terms of information contained in the Schwurbelarchiv, one substantial limitation of our dataset is the lack of unique user identifiers. As users can freely choose their usernames, the lack of access to user IDs makes it impossible to determine with certainty whether different authors are truly distinct individuals or whether posts with the same username originate from the same user. However, we believe the assumption that different usernames largely indicate activity by different individuals is warranted. To address this constraint, we plan to investigate text features and variance in temporal messaging patterns in the future to distinguish between authors who may share the same name.
We assessed the completeness of our dataset by comparing it to the content of comparable Telegram datasets [3, 13, 9]. This comparison also serves to estimate the scope of conspiracy-related discourse on Telegram during the COVID-19 pandemic in German speaking countries. On the level of chats the overlap with other datasets is low, with a maximum of 13% of the chats contained in the Schwurbelarchiv shared with Mohr [13]. This low overlap can largely be explained by a difference in inclusion criteria used to select chats, and differences in the temporal scope of the data sets. Another way to gauge the completeness of our dataset regarding conspiracy-related discourse is to look at message forwards. Here, we found that 54.1% of incoming forwarded messages are also contained as original messages in our dataset. On the flip side, this means that around half of the forwarded messages originate in chats that are not contained in our dataset. We therefore estimate that we cover around half of the conspiracy-related discourse happening in predominantly German-language Telegram chats at the time in our dataset. Other than potentially undersampling smaller chats from which fewer messages are forwarded, we have no reason to believe that the selection of chats included in our dataset is systematically biased. We therefore believe that the Schwurbelarchiv is an adequate resource for the study of conspiracy theories in the given temporal and geographic scope.
That being said, another limitation of our dataset relates to the unknown point in time at which data was collected for each chat. As has been reported by Buehling [15], messages on Telegram are deleted relatively frequently, and deletion patterns are not random. We estimate large portions of our dataset, particularly towards the end of the observation period, to have been collected via streaming and therefore be relatively unaffected by message deletion. However, we cannot determine the point at which streaming started for each chat. Nevertheless, we note that the other datasets covering a similar geographical and temporal scope have all been collected by bulk download through the Telegram API [9, 13, 3]. Therefore, these data sets are subject to increasing message deletion the farther a point in time is away from the time of data collection as well. By including both bulk-downloaded and streamed portions of chat content, the Schwurbelarchiv is likely more complete in terms of deleted messages than other data sets.
Next to the content of the messages, our data offers insights into the dynamics of information sharing and highlights how message frequency, group activity, and author engagement are distributed and evolve over time. The activity of authors in groups exhibits heavy-tailed distributions that approximate a power-law. This is consistent with activity patterns across social media platforms, such as Twitter [17]. Furthermore, activity distribution is very uneven, with the top 10% chats aggregating 73.3% of all messages and the top 10% authors contributing 94.6% of all messages.
The Schwurbelarchiv data can also be analysed from a network perspective by treating groups or authors as nodes and linking them according to shared membership or participation. This approach reveals topological features such as common affiliations or frequently co-occurring authors. Future research could harness these networks to investigate community structures, track the spread of (mis-)information and narratives, and identify influential participants across different discussions.
In summary, the Schwurbelarchiv provides broad utility for research in social network analysis, linguistic analysis, political science, and information studies, particularly in the realm of conspiracy theories. Future work can build on the dataset to investigate the dynamics of misinformation spread, conduct topic modelling, explore the influence of real-world events on user behaviour in online communities, and analyse the dynamics of conspiracy-related discourse in individual chats in a more qualitative way.
Declarations
[…]
Availability of data and materials The original Schwurbelarchiv dataset is available at the Internet Archive  [^11] https://github.com/cs2-lab/schwurbelarchiv. The cleaned and processed version of the dataset described in the paper is available upon reasonable request. Metadata of the dataset is available on Zenodo  [^12] https://zenodo.org/records/14704628. All code used to parse and analyze the data is publicly available on GitHub  [^13] https://github.com/cs2-lab/schwurbelarchiv.
[…]
Competing interests The author declare no competing interests.
[…]
Funding This research was funded in whole or in part by the Austrian Science Fund (FWF) [DOI:10.55776/P37280]. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. MA, EH & JP acknowledge the FWF funding.
[…]
Authors’ contributions MA, JL & JP conceived and designed the study. MA, EH, JL & JP performed data analysis and cleaning. MA, EH, JL & JP drafted the manuscript. All authors read and approved the final manuscript.
[…]
Acknowledgments The authors thank the anonymous individual who collected the original dataset, which made this work possible.
Appendix A Column description of the dataset
Column Description
uuid unique identifier for each row
folder_id internal folder identifier for the group/channel
chat_type type of chat (channel or group) based on number of different authors within a group (>3>3 authors == group); includes manual classification for 37 chat_names
chat_name name of the group/channel
author hashed name of author
link_url link as found in HTML a-tag
media_file name of attached media file
media_file_type voice message, video, or photo
message message content
message_id ID of message in respective group
posting_date date and time of message in the group (UTC-aware timestamptz)
replied_to message_id of message that has been replied to
website website domain, if link_url refers to a website
fwd_link_url link from forwarded message
fwd_media_file internal path of the forwarded media file
fwd_media_file_type voice message, video, or photo
fwd_message forwarded message content
fwd_posting_date date and time of message in the original group
message_lang language detected in message
is_conspiracy whether message was classified as conspiracy content
audio_transcription transcribed content of a speech-containing audio file
audio_length_s duration of an audio file (in seconds)
audio_size_mb file size of the attached audio file (in MB)
audio_lang language detected in audio transcription
video_transcription transcribed content of a speech-containing video file
video_length_s duration of a video file (in seconds)
video_lang language detected in video transcription
fwd_posting_date date and time of message in the original group (UTC-aware timestamptz)
fwd_author hashed name of the original author of the forwarded message
is_duplicate boolean flag marking rows that duplicate another due to redundant scraping of that group. Lines kept and flagged instead of deleted as for some semantic duplicates, the replied_to column references to only one of these lines.
Table 5: Schwurbelarchiv dataset columns.
Appendix B Base list of English keywords by [13]
keywords = { ’covid’, ’corona’, ’virus’, ’pandemic’, ’lockdown’, ’health’, ’mask’, ’distancing’, ’outbreak’, ’symptom’, ’quarantine’, ’influenza’, ’vaccine’, ’vaccination’, ’pandemic’, ’ventilator’, ’isolation’, ’immunity’, ’hospital’, ’icu’, ’intensive care unit’, ’treatment’, ’virologist’, ’clinic’, ’homeopathy’, ’sick’, ’pharma’, ’polymerase chain reaction’, ’pcr’, ’pertussis’, ’emergency’, ’injection’, ’doctors’, ’cellular’, ’remote work’, ’frontline’, ’covid-19’, ’sars’, ’sanitation’, ’pathogen’, ’propaganda’, ’disease’, ’epidemic’, ’diarrhea’, ’adjuvants’, ’respiratory’, ’hygiene’, ’protein’, ’medicine’, ’new cases’, ’positive’, ’scientist’, ’contagious’, ’mandate’, ’variant’, ’infect’, ’deaths’, ’ill’, ’cough’, ’measure’, ’viral’, ’mrna’, ’prevent’, ’healthcare’, ’contract’, ’shutdown’, ’smallpox’, ’booster’, ’antibody’, ’dose’, ’evidence’, ’misinformation’, ’isolation’, ’observed’, ’mandatory’, ’allergic’, ’allergy’, ’immune’, ’shortage’, ’syndrome’, ’drug’, ’chinese’, ’test’, ’restriction’, ’spread’, ’vax’, ’experimental’ }